OpenStudy (anonymous):

Suppose the area of a circle is 4pi. What is its circumference?

OpenStudy (anonymous):

Can you help me with a few more?

geerky42 (geerky42):

@micahm It is in reversed. s/he was given area and asked to find circumference, not other way around. Anyway, you can find radius from area using \(A=\pi r^2\), where \(A=4\pi\) Can you solve for r?

OpenStudy (anonymous):

BTW my school computer has blocked all websites but yahoo and openstudy.

OpenStudy (anonymous):

I'm a guy.

OpenStudy (anonymous):

r = 2 so d = 4 so the circumference is also 4pi

geerky42 (geerky42):

s/he = she or he. No need to tell your gender.

OpenStudy (anonymous):

So what if the area was 4pi and i was trying to find the radius?

geerky42 (geerky42):

Well, you first try to find radius from area, then you can plug in that value in formula of circumference \(C = 2\pi r\).

OpenStudy (anonymous):

I just got back from tearing through the house trying to catch my siblings. They had my calculator.

OpenStudy (anonymous):

@geerky42

geerky42 (geerky42):

Ok, so you know what radius is, right?

OpenStudy (anonymous):

Uh, yeah...

OpenStudy (isaiah.feynman):